OpenStudy (anonymous):

If the midpoint between (x, 7) and (10, 11) is (15, 9), find the value of x.

OpenStudy (anonymous):

You need to only look at the x's The midpoint is the average of the two x's (x + 10)/2 = 15 x + 10 = 30 x = 20 You can see the same is true for the y's: (7 + 11)/2 = 9 18/2 = 9 9 = 9